What is the volume of a cone with a radius of 4 centimeters and a height of 9 centimeters? Use 3.14 to estimate p.

Found 2 solutions by Boreal, Fombitz:
Answer by Boreal(15235)

V=(1/3)[pi*r^2*h]=(1/3)(pi*16*9)=48 pi cm^3=150.72 cm^3.

Answer by Fombitz(32388)